OpenStudy (anonymous):

Will someone please explain to me how to work these differentiation logarithmic functions? :(

OpenStudy (shadowfiend):

What particular issues are you having?

OpenStudy (shadowfiend):

i.e., do you have a specific problem you can show that you're having trouble with?

OpenStudy (anonymous):

Sure :D \[dy/dx=x/(1+x^2) \] and I'm supposed to solve that by separating and integrating i believe, using the ln and then "e"-ing both sides to get rid of it, but I don't understand how!! Thank you! :)

OpenStudy (shadowfiend):

Hm. Ok, I don't have time to do the entire thing, but basically the starting point is the fact that using ln you can turn a division into a subtraction. So: \[\ln \frac{a}{b} = \ln a - \ln b\] You can use that to do: \[\ln \frac{dy}{dx} = \ln \frac{x}{1 + x^2}\] \[\ln \frac{dy}{dx} = \ln x - \ln(1 + x^2)\] Then you can integrate the two sides. Does that help?

OpenStudy (anonymous):

Yes (: Thank you

OpenStudy (shadowfiend):

Awesome! Good to hear :)